matter has a body-centered cubic unit cell with a length of 330 π‘π‘š along an edge


and a density of 8.57 𝑔 π‘π‘šβˆ’3


. Calculate the atomic mass of the matter.

1
Expert's answer
2021-12-09T10:09:20-0500

The atomic mass:


"a^3=\frac{2M}{\rho N_A},\\\space\\
M=a^3\rho N_A/2=9.3·10^{-5}\text{ g/mol}."